OpenStudy (anonymous):

I'm really confused right now Okay so my problem tells me to write a function rule for the area of a triangle whose base is 4 feet more than its height. The choices are : A. A = 0.5h2 + 4h; 42 ft2 B. A = 4h − 0.5h2; 6 ft2 C. A = 0.5h2 + 2h; 30 ft2 D. A = 0.5h2 − 2h; 6 ft2 Please explain and not just give me the answer. Believe it or not I actually don't want to fail Algebra.

OpenStudy (bibby):

\(area=\frac{1}{2}b*h\) if b=h+4 replace it and you'll get \(a=\frac{1}{2}(h)(h+4)=\frac{1}{2}(h^2+4h)=\frac{h^2}{2}*2h\)

jimthompson5910 (jim_thompson5910):

on the last part, bibby probably meant to say \(\Large \frac{h^2}{2} + 2h\) and not \(\Large \frac{h^2}{2} * 2h\)

jimthompson5910 (jim_thompson5910):

keep in mind that \[\Large \frac{h^2}{2} = \frac{1}{2}*h^2 = 0.5h^2\] since 1/2 = 0.5

OpenStudy (bibby):

I'm not sure why there's a part 2 that says 30 but yeah I meant \(0.5h^2+2h\)

jimthompson5910 (jim_thompson5910):

I'm guessing something got cut off like "Part 2: Find the area when h = ... "

OpenStudy (bibby):

yes

OpenStudy (bibby):

0.5h^2+2h=30 0.5(h^2+4h)=30 h^2+4h=60 h^2+4h-60=0 (h-6)(h+10)=0 h=6

OpenStudy (anonymous):

hey guys, sorry i didn't know someone would answer so quick. i think i understand so thanks . I just put part 1 up cuz i thought that if someone could explain Part 1 i could do Part2, so thank you ! Bibby and Jim r the best!

OpenStudy (bibby):

well once they give you h=6 you just plug it into the equation we found in part 1. 0.5*6*6+2*6 0.5*36+12 18+12=30